OpenStudy (anonymous):

Find the standard equation of the circle passing thru a given point with a given center:: center (1,5) passing thru (3,2)

OpenStudy (anonymous):

@LindseyLou877 Do you know circle formula?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Please find it from your note or textbook, then I'll show you how to plug the numbers in!

OpenStudy (anonymous):

(x-h)^2+(y-k)^2=r^2

OpenStudy (anonymous):

?

OpenStudy (anonymous):

im taking a timed quiz and cant figure ths one out

OpenStudy (anonymous):

Great :) Try to memorize it! Center ( h, k) = ...?

OpenStudy (anonymous):

1 and 5

OpenStudy (anonymous):

Now, plug the center into (x -h)² + ( y-k)² = ... ?

OpenStudy (anonymous):

I can get that far.. this is where I am having the problem and I am running outta time

OpenStudy (anonymous):

Plug 1 into where h is, similarly plug 5 ....

OpenStudy (anonymous):

and thats my answer?seems too simple

OpenStudy (anonymous):

Follow my instruction, plug them in and show me on HERE!

OpenStudy (anonymous):

(X-1)^2+(y-5)^2=

OpenStudy (anonymous):

r^2

OpenStudy (anonymous):

Which is the distance! Do you know distance formula?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I have to hurry in order to get this in time

OpenStudy (anonymous):

I don't like short "no", please find the formula yourself if you want to learn!

OpenStudy (anonymous):

I get sq rt 13 when i used distance formula

OpenStudy (anonymous):

Correct, now R= 13 Plug it into the right side, that's it!

OpenStudy (anonymous):

I am having a probem figuring out how to write this out

OpenStudy (anonymous):

(x-1)^2+(y-5)^2=13

OpenStudy (anonymous):

or (x-1)² + ( y -5)² = (√ 13)²

OpenStudy (anonymous):

do i need the sq rt sign in my answer or not ?

OpenStudy (anonymous):

I hope you memorize the formula from now on, because it's plain simple to plug the data in! Keep in mind that you'll fail the test by saying "no"!

OpenStudy (anonymous):

It's better to keep (√ 13)², as in the formula r²

OpenStudy (anonymous):

it says to simplify thou so

OpenStudy (anonymous):

Either way is fine, because (√ 13)² = 13, the same result!

OpenStudy (anonymous):

Of course I know the formulas, but what the point because YOU are the one will taking the test, not me!